the distance between  the centres of carbon and oxygen atoms in the carbon monoxide gas molecule is 1.13 * 10' 10 m . the distance of centre of mass of the molecule relative to carbon atom is

Grade:11

1 Answers

Manas Shukla
7 years ago
Let the mass of carbon be 12 units and oxygen be 16 units (12 u , 16 u),
let A = distance between the two.
Lets assume carbon atom be at origin, So,
centre of mass relative to carbon = \frac{12\times 0 +16\times A}{16+12}
 = 4A/7 = 1.13\times 10^{-10} \times \frac{4}{7}
=6.46 \times 10^{-11}m